40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 13 03:12:47 2013, Manasa, Jenne, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 08:34:33 2013, yuta, Update, Locking, TWO arms TWO colors 
       Reply  Wed Mar 13 16:29:25 2013, Manasa, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 10:26:58 2013, Manasa, Update, Locking, We're still good!! 
    Reply  Mon Mar 25 15:01:33 2013, Manasa, Update, Locking, TT1 drift 
       Reply  Mon Mar 25 15:24:41 2013, rana, Update, Locking, TT1 drift 
Message ID: 8284     Entry time: Wed Mar 13 10:26:58 2013     In reply to: 8282
Author: Manasa 
Type: Update 
Category: Locking 
Subject: We're still good!! 

We're still good with the IFO alignment after 7hours.

I found the green still locked in the same state as last night; but no IR (so the arms are stable and the TTs should definitely take the blame).

From last night's observation (elog about drift in TT1), I only moved TT1 in pitch and gained back locking in IR for both the arms

ELOG V3.1.3-